Differentiate the function 𝑓(π‘₯) = (3π‘₯ + 1) 4 (2π‘₯ βˆ’ 1) 5 and simplify your answer


1
Expert's answer
2022-03-30T08:46:43-0400

"f(x)=(3x+1)\u2074(2x-1)\u2075"


Let "u=(3x+1)⁴ =>\frac{du}{dx}=12(3x+1)³"


"v=(2x-1)⁵=>\frac{dv}{dx}=10(2x-1)⁴"


"f'(x)=v\frac{du}{dx}+u\frac{dv}{dx}"


"f'(x)=(2x-1)\u2075\u202212(3x+1)\u00b3+(3x+1)\u2074\u202210(2x-1)\u2074"


"f'(x)=2(2x-1)\u2074(3x+1)\u00b3[6(2x-1)+5(3x+1)]"


"f'(x)=2(2x-1)\u2074(3x+1)\u00b3[12x-6+15x+5]"


"f'(x)=2(2x-1)\u2074(3x+1)\u00b3[27x-1]"